OpenStudy (supernova_sonntag):

Solve algebraically: y=x^2-20 y=x-8 I have tried this many times and I can't figure it out.

OpenStudy (pawanyadav):

Plug value of y from eq 1 to equation 2, We get x^2-20=x-8 x^2-x-20+8=0 x^2-x-12=0 x^2-4x+3x-12=0 x(x-4)+3(x-4)=0 (x-4)(x+3)=0 Can you solve now.

OpenStudy (pawanyadav):

@Supernova_Sonntag

OpenStudy (anonymous):

it is 4

OpenStudy (pawanyadav):

4 and -3.(both)

OpenStudy (supernova_sonntag):

Yes that is all I needed, thank you. I guess I was getting hung up on the factoring.

OpenStudy (pawanyadav):

You are welcome